OpenStudy (anonymous):

find the perimeter of a rectagular farm that is 1/2 mile wide and 3/4 mile long.

OpenStudy (anonymous):

P = 2(L+B)

OpenStudy (ghazi):

perimeter of rectangle =2(length +breadth)

OpenStudy (anonymous):

= 2* (1/2 + 3/4) = 2* 5/4 = 2.5